# Large cluster formation in a simple model for globular colloidal aggregates

### v2

@@ II Model and methods
Molecular dynamics simulations have been run specifically in the NVT ensemble using the LAMMPS packageThompson et al. (2021) for a total of 218000 particles of mass 100 amu. Distance and energy units for the simulation have been chosen following the LAMMPS package real units specification. The integration time step was set to 2 fs. Results will be presented in terms of reduced units, both for temperature (defined in terms of the well depth as indicated above), length (using $\sigma=4$Å  as unit length) and time, for which the time unit is given by $\tau_{0}=(m\sigma^{2}/\zeta)^{1/2}=7.5$ps. The systems were equilibrated for 20 ns, of which the first 10 ns correspond to a slow cooling process following a linear temperature ramp starting from a fully disordered system at kBT/$\zeta=18$ down to the desired temperatures, namely, kBT/$\zeta$ = 3, 6, 9, 12, and 15 for $\rho\sigma^{3}=0.05$ and just kBT/$\zeta$ = 3, and 9 for $\rho\sigma^{3}=0.025$. It is worth stressing that shorter equilibration runs (5 ns) lead to very similar cluster size/radii distributions. This is an indication that the results presented correspond to equilibrium states. For much shorter equilibration runs and/or cooling times it is very easy to end up with bimodal or even multi-modal cluster size/radii distributions. Production runs were $\tau_{run}=20$ns long with configurations stored every 10 ps.
@@
Figure 3: Average life time of the clusters, ($\tau_{av}$, red bars) as a function of temperature for $\rho\sigma^{3}=0.05$ (blue bars) and $\rho\sigma^{3}=0.025$ (red bars). The upper limit of the $y$-axis corresponds to the length of the simulation production run, $\tau_{run}$.
@@
Figure 4: a) Cluster size distribution dependence on temperature and global density b) Cluster radii of gyration distribution in terms of density and temperature. Average radii are indicated by vertical lines.
@@ III Onset of clustering
As mentioned before, we have chosen potential parameters and thermodynamic conditions for the system to yield a globular cluster phase. We have first explored a series of temperatures for the two densities indicated above using the Hypernetted Chain integral equation Hansen and McDonald (2006). As shown by Archer and WildingArcher and Wilding (2007) this equation has a non-solution region which hides the cluster phase. This lack of solutions (in fact solutions have analytical continuation in the complex spaceLomba and López-Martín (1995)) can be interpreted as a signature of a phase transition, in this instance a first order transition between the particle gas and a fluid of liquid-like spherical clusters. In our case we find that real solutions for the largest density disappear below $k_{B}T_{ns}^{HNC}/\zeta=14.9$, and $k_{B}T/\zeta=8$ for $\rho\sigma^{3}=0.025$. For $k_{B}T_{c}/\zeta=17.5$ and $\rho\sigma^{3}=0.05$ the HNC converges yielding a structure factor in excellent agreement with the simulation (cf. Figure 2). A visual inspection of simulation snapshots does not show evidence of clustering. Interestingly, in Figure 2 we first see that for this temperature there is a significantly high prepeak (also known as intermediate range order peak Godfrin et al. (2014)) which reaches the value $S(Q_{m})\approx 2.7$ for $Q_{m}\approx 2\pi/\lambda_{0}$, whereas for $Q\sigma\gtrsim 1$ the structure factor corresponds to that of a dilute gas of uncorrelated particles. Godfrin and coworkers propose the use of this precise value of the prepeak height as signature of the onset of the cluster phaseGodfrin et al. (2014). We will see in the next sections that lowering the temperature increases the height of the prepeak by orders of magnitude, i.e. $S(Q_{m})$ practically diverges for a $Q_{m}\neq 0$, which suggests the existence of a non-zero Lifshitz pointBomont and Costa (2012). This means that the system becomes unstable with respect to a modulated inhomogeneous phase, whose modulation is given by $Q_{m}\approx 2\pi/\lambda_{0}$. The locus of Lifshitz points for varying densities yields the $\lambda$-line Archer and Wilding (2007), a term borrowed from the study of criticality in ionic fluidsStell (1995). It must be mentioned, that Bollinger and coworkersBollinger and Truskett (2016) suggested the use of an additional criterion to characterize the onset of the cluster phase, namely, a thermal correlation length with values within the interval $2\lesssim\xi_{T}/\sigma\lesssim 3$. This quantity can be estimated from a Lorentzian fit of the prepeak, which in our case yields $\xi_{T}/\sigma=5.6$, exceeding the limit suggested in Ref. Bollinger and Truskett (2016). However, one must bear in mind that Ref. Bollinger and Truskett (2016) is focused on Yukawa (screened Coulomb) interactions, and here we are dealing with bare exponentials. So far it has not been assessed to what extent the choice of the explicit form of the potential affects the reliability of this criterion. Until a complete analysis of the type performed in Ref. Bollinger and Truskett (2016) is carried out it seems reasonable to stick to Godfrin’s et al choice.
@@ IV Single-cluster structure and dynamics
In order to define the clusters we have used a geometric criterion, namely, all those particles whose separation is below $d_{cl}=1.3\sigma$ are considered as pertaining to the same cluster. This distance roughly corresponds to the first non-zero minimum of the particle-particle pair distribution function, i.e. it is an estimate of the outer boundary of the first coordination shell. Once this clustering distance was set, we have run a GPU density based parallel scan (G-DBSCAN Andrade et al. (2013)) using an in-house code. Specific GPU optimizations were performed to calculate structure factors, pair distribution functions Levine et al. (2011); Sakharnykh (2015); Reuter and Köfinger (2019) and dynamic properties. Only clusters containing more than 4 particles have been considered for the analysis below.
@@
A first observation from our cluster analysis is the fact that the structures are long-lived. In Figure 3 the average cluster life time, $\tau_{av}$ is displayed for the thermodynamics states under consideration. The top of the ordinate axis corresponds to the length of the production run, $\tau_{run}\approx 2700\tau_{0}$. For the highest density, up until $k_{B}T/\zeta=15$ the average lifetime exceeds half the length of the simulation, which is an indication of the stability of the structures. This is explained by the large depth of the cluster-particle potential (see right graph of Figure 1). At $k_{B}T_{c}/\zeta\approx 15$ the clusters are extremely short lived, this temperature being fairly close to the limit of real solutions of the Hypernetted Chain integral equation, which has been discussed above. For higher temperatures no significant clustering is found. Now, when the density decreases the average life time of the clusters decreases substantially. The real solution of the HNC disappears at $k_{B}T/\zeta=7.6$, but for $k_{b}T/\zeta=9$ the simulation results still display a significant degree of clustering, although much more short lived. Simulation runs for higher temperatures (results not included) do not show evidence of clusters satisfying our minimum size criterion.
@@
Figure 5: Temperature and density dependence of the average density profile of the clusters. Densities have been scaled with the effective diameter of the particles which is $\sigma_{eff}\approx 0.8\sigma$. Vertical lines denote the radii corresponding to the Gibbs dividing surface (cf. Eq. (6) in the text).
@@
Figure 6: Upper graph) Single particle mean square displacement dependence of temperature and density. Lower graph) Density and temperature dependence of the center of mass mean square displacement for persistent clusters. In both graphs velocities and time are expressed in reduced units.
@@
$R_{g}=\left<\sqrt{\frac{\sum_{i}m_{i}({\bf r}_{i}-{\bf R}_{com})^{2}}{\sum_{i}m_{i}^{2}}}\right>$ (5)
@@
where the brackets indicate a ensemble average and ${\bf R}_{com}$ is the center of mass of the cluster whose radius of gyration is being computed. We see that for the largest density the size distribution peaks at around 450 particles and at 200 particles for the lowest. For all temperatures the clusters are liquid droplets, as will be illustrated below when analyzing the short time behavior of the mean square displacement.
@@
Interestingly one can see that the shape of the cluster size distribution is hardly affected by temperature in the case of the largest density, whereas for low density becomes much wider and much larger clusters appear as temperature rises. This, in combination with the short lifetime of the clusters seen in Figure 3, is an indication of a rapid merging and splitting of clusters taking place for the lower density/high T case. The large spread of the low density-high T size distribution is also reflected in the cluster radii distribution, which in this case is almost Gaussian-like. All remaining distributions are negatively skewed beta-distributions with a more or less wide hump at approximately one particle diameter below the maximum of $P(R_{g})$. This probably results from the tendency of clusters to free and recapture their external layer of particles. This is more difficult for internal layers where particles are strongly attached, except for the low density/high temperature case. Here dominates the thermal equilibrium between destruction and creation of clusters leading to the Gaussian-like distribution.
@@
This picture is complemented by an inspection of the cluster average density profiles, $\rho_{cl}(r)$, depicted in Figure 5. In this case, we have scaled the densities with $\sigma_{eff}=0.8\sigma$ in order to be able to compare the densities with those of standard fluids (e.g. for a Lennard-Jones fluid, liquid densities are usually assumed to be $\rho\sigma^{3}\gtrsim 0.4$). One first notices that despite the extremely low total particle densities, for $\rho\sigma^{3}=0.05$ the clusters display very high values of the effective density profile, even for high T. These values are well into the liquid domain . For the lower global density, the effective inner cluster density remains almost within the boundaries of the liquid state. In both cases there is a transition towards gas-like densities as the profile approaches the surface in a gradual fashion. Taking into account that the clusters are rather spherical, we have defined a Gibbs dividing surface in order to have a quantitative description of the separation between gas-like particles and liquid-like particles, as is customary in the description of gas-liquid interfaces Hansen and McDonald (2006). This quantity places an ideal separation between the gas-like cluster region from its corresponding liquid-like counterpart. Here, for a spherical cluster this should be the surface of a sphere, whose radius (Gibbs dividing radius) $R_{Gibbs}$ is given by
@@
We see that this Gibbs radii separating gas-like and liquid-like particles in the clusters (vertical lines in Figure 5) follow the same trends as the average radii of gyration (vertical lines in Figure 4), i.e. larger values for larger densities, and for the same density the radius increases with temperature. This latter trend reflects the larger spread of the cluster size distributions as temperature raises. The Gibbs radii are systematically larger ($\approx 0.5\sigma$) than the gyration radii. This is a consequence of the lesser weight of the “gas atmosphere” surrounding the quasi-spherical clusters in Eq. (5). It is worth noticing that the cluster density profiles reach vanishing values in the interval $4.5\sigma<r<5.5\sigma$, by which the average cluster spatial extent is $\lesssim\lambda_{0}$.
@@
Figure 7: Vibrational frequency spectra derived from the particles velocity self-correlation at various temperatures and $\rho\sigma^{3}=0.05$, and the cluster vibrational frequency spectrum (inset in lower graph) as computed from the velocity self-correlation of the persistent clusters centers of mass.
@@
Figure 8: Cluster-cluster pair distribution function and intra-cluster pair distribution function (inset) for the lowest temperature state $k_{B}T/\zeta=0.3$.
@@
Now a few words concerning the single particle dynamics. In what follows, in order to make the results independent of specific details of the model, time and frequency will be reduced with the corresponding molecular time unit, $\tau_{0}$ and its inverse $\tau_{0}^{-1}$ respectively. Here we first look at the individual particle mean square displacement (m.s.d), plotted in the upper graph of Figure 6. It is apparent from the figure that there are clearly two separate regimes. First, a short time scale (under 10 $\tau_{0}$, i.e. for our system around 70ps) where the system is in a diffusive regime corresponding to intracluster movements. This regime can be split into two more or less linear regions with a decreasing diffusion constant (slope of the curves). This reflects the finite size of the clusters. For much longer times (up to 2.7×$\times 10^{3}\tau_{0}$) the behavior of the mean square displacement is represented in the inset. We observe for the low density-high temperature case a fully linear behavior, with the m.s.d. exhibiting the steepest slope, i.e. the largest diffusivity. Except for the high density-low temperature case the two remaining cases display a plateau after the initial diffusive regime followed by a long time diffusive behavior. In this latter region the mean square displacement grows rather linearly as well. This is contrast with the higher density-low temperature case in which the plateau extends for the duration of the simulation. This long time behavior stems mostly from the collective cluster movement which is shown in the lower graph of Figure 6. Here we have plotted the mean square displacement of the center of mass of persistent clusters, that is, those that preserve their identity during the length of the production run, neither splitting nor merging with other clusters. At short times one can even appreciate the ballistic regime with a parabolic m.s.d. before the clusters begin to collide. We see that the long time behavior (inset) displays the same trends as the single particle mean square displacement, but one can actually measure a tiny diffusivity for the clusters at higher density/low temperature, namely $D/(\sigma^{2}\zeta/m)^{1/2}=0.0003$. This is an indication that our system is approaching an arrested glassy state. Taking into account that the source of this quasi-freezing of the cluster positions is mostly induced by the long range of the repulsive interactions (cf the right graph of Fig. 1) one is tempted to identify this state as a precursor of a Wigner glass of clustersRuiz-Franco and Zaccarelli (2021). However, strictly speaking the formation of Wigner crystals and glasses Duang and Guojun (2005) is due to extremely long ranged interactions such a Coulombic ones, in conjunction with higher packing fractions. Following the discussion of Klix and coworkers Klix et al. (2010) and bearing in mind the significant interparticle attraction of our model and low packing fraction, we are probably somewhere near the dynamic transition towards a cluster glassy state.
These features are further confirmed by an analysis of the velocity self-correlation functions (not represented for the sake of brevity). In the case of the low T-high density clusters there is a well separated time regime in which the particles within the cluster experience the “cage effect”, i.e. density is sufficiently high for the particles to reverse direction due to collisions with the nearest neighbors. This is less apparent as temperature increases and density is lowered. These effects are reflected in the vibrational frequency spectrum, which is the Fourier transform of the particle velocity self-correlation function in the time domain. This quantity is depicted in Figure 7 for low and high temperatures and the largest density, both for the single particle dynamics and in the inset for that of the centers of mass of persistent clusters, $Z_{cl}(\nu)$. Thus, concerning particle diffusivities, these correspond to the values of $Z(0)$ ( with our convention for the velocity autocorrelation $Z(0)=D$). In agreement with the behavior of the m.s.d. (cf. Figure 6), diffusivity increases with temperature, and we see that from the point of view of the particle dynamics the system is fluid-like. At the lowest temperature we observe a small secondary maximum close to $\nu\approx 10^{-12}\tau_{0}^{-1}$ (cf. lower inset in Figure 7): this corresponds to “vibrational” movements due to particle collisions inside the clusters, i.e. the aforementioned “cage” effect. The first maximum that occurs at $\nu\approx 0.5\times 10^{-12}\tau_{0}^{-1}$ for the highest temperature is connected with traslational excitations, and it shifts to lower frequencies as temperature is decreased. .
@@
Figure 9: Cluster-cluster structure factor for the lowest temperature state $k_{B}T/\zeta=0.3$.
In the inset we present the vibrational spectra obtained from the cluster’s center-of-mass velocity self-correlation. It displays very definite peaks at substantially low frequencies. These peaks correspond to collective cluster vibrations around the equilibrium positions that are almost frozen into an arrested state, as can be inferred from their relatively low diffusivities, both computed from the long time limit of the m.s.d. and from $Z_{cl}(0)$ in Figure 7. The effect of temperature for the density in question on the cluster dynamics is precisely to increase the intensity of the collective vibrations of the clusters. Obviously, for $k_{B}T/\zeta>15$ this spectrum vanishes due to the dissociation of the aggregates and subsequent melting of the cluster fluid into a low density homogeneous fluid.
V Two-particle and intercluster structural correlations

### v4

@@ II Model and methods
Molecular dynamics simulations have been run specifically in the NVT ensemble using the LAMMPS packageThompson et al. (2021) for a total of 218000 particles of mass 100 amu. Distance and energy units for the simulation have been chosen following the LAMMPS package real units specification. The integration time step was set to 2 fs. Results will be presented in terms of reduced units, both for temperature (defined in terms of the well depth as indicated above), length (using $\sigma=4$Å  as unit length) and time, for which the reduced time unit is given by $\tau_{0}=(m\sigma^{2}/\zeta)^{1/2}=7.5$ps. The systems were equilibrated for 20 ns, of which the first 10 ns correspond to a slow cooling process following a linear temperature ramp starting from a fully disordered system at kBT/$\zeta=18$ down to the desired temperatures, namely, kBT/$\zeta$ = 3, 6, 9, 12, and 15 for $\rho\sigma^{3}=0.05$ and just kBT/$\zeta$ = 3, and 9 for $\rho\sigma^{3}=0.025$. It is worth stressing that shorter equilibration runs (5 ns) lead to very similar cluster size/radii distributions. This is an indication that the results presented correspond to equilibrium states. For much shorter equilibration runs and/or cooling times it is very easy to end up with bimodal or even multi-modal cluster size/radii distributions. Production runs were $\tau_{run}=20$ns long with configurations stored every 10 ps.
@@
Figure 4: Average life time of the clusters, ($\tau_{av}$, as a function of temperature for $\rho\sigma^{3}=0.05$ (blue bars) and $\rho\sigma^{3}=0.025$ (red bars). The upper limit of the $y$-axis corresponds to the length of the simulation production run, $\tau_{run}$. Note that for $\rho\sigma^{3}=0.025$ the only temperatures calculated are $k_{B}T/\zeta=3,$ and 9.
@@
Figure 5: a) Cluster size distribution dependence on temperature and global density b) Cluster radii of gyration distribution in terms of density and temperature. Average radii are indicated by vertical lines, with solid ones corresponding to $\rho\sigma^{3}=0.05$ and dash-dotted ones to $\rho\sigma^{3}=0.025$.
@@ III Onset of clustering
As mentioned before, we have chosen potential parameters and thermodynamic conditions for the system to yield a globular cluster phase. We have first explored a series of temperatures for the two densities indicated above using the Hypernetted Chain integral equation Hansen and McDonald (2006). As shown by Archer and WildingArcher and Wilding (2007) this equation has a non-solution region which hides the cluster phase. This lack of solutions (in fact solutions have analytical continuation in complex spaceLomba and López-Martín (1995)) can be interpreted as a signature of a phase transition, in this instance a first order transition between the particle gas and a fluid of liquid-like spherical clusters. For a very closely related system, this first order transition was identified by Archer and WildingArcher and Wilding (2007) using both Density Functional Theory and extensive Grand Canonical Monte Carlo simulations exploiting multi-canonical reweighting aided by multi-histogram reweighting techniquesArcher and Wilding (2007). In our case we find that real solutions for the largest density disappear below $k_{B}T_{ns}^{HNC}/\zeta=14.9$, and $k_{B}T/\zeta=8$ for $\rho\sigma^{3}=0.025$. For $k_{B}T/\zeta=17.5$ and $\rho\sigma^{3}=0.05$ the HNC converges yielding a structure factor in excellent agreement with the simulation (cf. Figure 3). A visual inspection of simulation snapshots does not show evidence of clustering. Interestingly, in Figure 3 we first see that there is a significantly high prepeak (also known as intermediate range order peak Godfrin et al. (2014)) which reaches the value $S(Q_{max})\approx 2.7$ for $Q_{max}\approx 2\pi/\lambda_{0}$, whereas for $Q\sigma\gtrsim 1$ the structure factor corresponds to that of a dilute gas of uncorrelated particles (cf. the inset in Figure 3). Godfrin and coworkers propose the use of this precise value of the prepeak height as signature of the onset of the cluster phaseGodfrin et al. (2014). We will see in the next sections that lowering the temperature increases the height of the prepeak by orders of magnitude, i.e. $S(Q_{max})$ practically diverges for a $Q_{max}\sim Q_{0}\neq 0$, which suggests the existence of a non-zero Lifshitz pointBomont and Costa (2012). This means that the system becomes unstable with respect to a modulated inhomogeneous phase, whose modulation is given by $Q_{max}\approx 2\pi/\lambda_{0}$. The locus of Lifshitz points for varying densities yields the $\lambda$-line Archer and Wilding (2007), a term borrowed from the study of criticality in ionic fluidsStell (1995). It must be mentioned, that Bollinger and coworkersBollinger and Truskett (2016) suggested the use of an additional criterion to characterize the onset of the cluster phase, namely, a thermal correlation length with values within the interval $2\lesssim\xi_{T}/\sigma\lesssim 3$. This quantity can be estimated from a Lorentzian fit of the prepeak, which in our case yields $\xi_{T}/\sigma=5.6$, exceeding the limit suggested in Ref. Bollinger and Truskett (2016). However, one must bear in mind that Ref. Bollinger and Truskett (2016) is focused on Yukawa (screened Coulomb) interactions, and here we are dealing with bare exponentials. So far it has not been assessed to what extent the choice of the explicit form of the potential affects the reliability of this criterion. Until a complete analysis of the type performed in Ref. Bollinger and Truskett (2016) is carried out it seems reasonable to stick to Godfrin’s et al choice.
An illustration of a globular cluster configuration can be seen in Figure 2, where we have also represented the effective excluded volume of the clusters using transparent spheres. One can see that despite a relatively low concentration of clusters, the spheres representing the excluded volume are arranged in a rather dense packing, which will be reflected both in the microscopic structure and dynamics of our system.
@@ IV Single-cluster structure and dynamics
In order to define the clusters we have used a geometric criterion, namely, all those particles whose separation is below $d_{cl}=1.3\sigma$ are considered as pertaining to the same cluster. This value is obtained from a preliminary run at the lowest temperature and $\rho\sigma^{3}=0.05$ for which the particle-particle distribution function, $g(r)$, was evaluated. We found that $d_{cl}=1.3\sigma$ roughly corresponds to the first non-zero minimum of $g(r)$, i.e. it is an estimate of the outer boundary of the first coordination shell. Once this clustering distance was set, we have run a GPU density based parallel scan (G-DBSCAN Andrade et al. (2013)) using an in-house code. Specific GPU optimizations were performed to calculate structure factors, pair distribution functions Levine et al. (2011); Sakharnykh (2015); Reuter and Köfinger (2019) and dynamic properties. Only clusters containing more than 4 particles have been considered for the analysis below.
@@
A first observation from our cluster analysis is the fact that the structures are long-lived. In Figure 4 the average cluster life time, $\tau_{av}$ is displayed for the thermodynamics states under consideration. This quantity is defined as the average of the time a given cluster retains its identity (does not merge, split or dissolve, but it might incorporate/lose individual particles) in the course of the simulation. The top of the ordinate axis corresponds to the length of the production run, $\tau_{run}\approx 2700\tau_{0}$. For the highest density, up until $k_{B}T/\zeta=15$ the average lifetime exceeds half the length of the simulation, which is an indication of the stability of the structures. This is explained by the large depth of the cluster-particle potential (see right graph of Figure 1). At $k_{B}T_{c}/\zeta\approx 15$ the clusters are extremely short lived, this temperature being fairly close to the limit of real solutions of the Hypernetted Chain integral equation, which has been discussed above. For higher temperatures no significant clustering is found. Now, when the density decreases the average life time of the clusters decreases substantially. The real solution of the HNC disappears at $k_{B}T/\zeta=7.6$, but for $k_{b}T/\zeta=9$ the simulation results still display a significant degree of clustering, although much more short lived. Simulation runs for higher temperatures (results not included) do not show evidence of clusters satisfying our minimum size criterion. In Figures S1 in the Supplementary Information one can see snapshots of the evolution of the cluster formation as the system is cooled down from $k_{B}T/\zeta=18$ to $k_{B}T/\zeta=3$.
@@
Figure 6: Temperature and density dependence of the average density profile of the clusters. Densities have been scaled with the effective diameter of the particles which is $\sigma_{eff}\approx 0.8\sigma$. Vertical lines denote the radii corresponding to the Gibbs dividing surface (cf. Eq. (8) in the text), with solid ones corresponding to $\rho\sigma^{3}=0.05$ and dash-dotted ones to $\rho\sigma^{3}=0.025$.
@@
Figure 7: (Upper graph) Single particle mean square displacement dependence on temperature and density. (Lower graph) Density and temperature dependence of the center of mass mean square displacement for persistent clusters. In both graphs velocities and time are expressed in reduced units. In the insets the long time behavior is illustrated using log-log plots.
@@
$R_{g}=\left<\sqrt{\frac{\sum_{i}m_{i}({\bf r}_{i}-{\bf R}_{com})^{2}}{\sum_{i}m_{i}}}\right>$ (7)
@@
where the brackets indicate a ensemble average and ${\bf R}_{com}$ is the center of mass of the cluster whose radius of gyration is being computed. We see that for the largest density the size distribution peaks at around 450 particles and at 200 particles for the lowest. For all temperatures the clusters are liquid droplets, as will be shown below when analyzing the short time behavior of the mean square displacement.
@@
Interestingly one can see that the shape of the cluster size distribution is hardly affected by temperature in the case of the largest density, whereas for low density it becomes much wider and much larger clusters appear as temperature rises. This, in combination with the short lifetime of the clusters seen in Figure 4, is an indication of a rapid merging and splitting of clusters taking place for the lower density/high T case. The large spread of the low density-high T size distribution is also reflected in the cluster radii distribution, which in this case is almost Gaussian-like. All remaining distributions are negatively skewed beta-distributions with a more or less wide hump at approximately one particle diameter below the maximum of $P(R_{g})$. Large clusters are in equilibrium with smaller ones, coalescing with them and letting particles escape their outer atmosphere to subsequently form smaller clusters. Merging of larger clusters is prevented by the large intercluster repulsion (cf. see right graph of Figure 1), and this is at the root of the beta-distribution shapes. For the low density/high temperature case thermal energy is sufficient to overcome the intercluster repulsion to a certain extent, and the equilibrium between destruction and creation of clusters leads to a Gaussian-like distribution.
@@
This picture is complemented by an inspection of the cluster average density profiles, $\rho_{cl}(r)$, depicted in Figure 6. In this case, we have scaled the densities with $\sigma_{eff}=0.8\sigma$ in order to be able to compare the densities with those of standard fluids (e.g. for a Lennard-Jones fluid, liquid densities are usually assumed to be $\rho\sigma^{3}\gtrsim 0.4$). One first notices that despite the extremely low total particle densities, for $\rho\sigma^{3}=0.05$ the clusters display very high values of the effective density profile, even for high T. These values are well into the liquid domain. For the lower global density, the effective inner cluster density remains practically within the boundaries of the liquid state. In both cases there is a transition towards gas-like densities as the profile approaches the surface in a gradual fashion. Taking into account that the clusters are rather spherical, we have defined a Gibbs dividing surface in order to have a quantitative description of the separation between gas-like particles and liquid-like particles, as is customary in the description of gas-liquid interfaces Hansen and McDonald (2006). This surface places an ideal separation between the gas-like cluster region from its corresponding liquid-like counterpart. Here, for a spherical cluster this should be the surface of a sphere, whose radius (Gibbs dividing radius) $R_{Gibbs}$ is given by
@@
We see that these Gibbs radii separating gas-like and liquid-like particles in the clusters (vertical lines in Figure 6) follow the same trends as the average radii of gyration (vertical lines in Figure 5), i.e. larger values for larger densities, and for the same density the radius increases with temperature. This latter trend reflects the larger spread of the cluster size distributions as temperature raises. The Gibbs radii are systematically larger ($\approx 0.5\sigma$) than the gyration radii. This is a consequence of the lesser weight of the “gas atmosphere” surrounding the quasi-spherical clusters in Eq. (7). It is worth noticing that the cluster density profiles reach vanishing values in the interval $4.5\sigma<r<5.5\sigma$, by which the average cluster spatial extent is $\approx\lambda_{0}/2$, in accordance with the estimated obtained from the potential’s characteristic wave vector (cf. Figure 1). The temperature and density dependence of both $R_{Gibbs}$ and $R_{g}$ found in this paper agrees with the finding of Schwanzer et al. Schwanzer et al. (2016) for much smaller clusters in two dimensions. It is important to stress that in our system further increases of density maintaining the temperature fixed would lead to an increase in the number and size of the clusters, that would eventually coalesce to form lamellar phases which finally mutate into a bubble phase. Other SALR interactions (or different potential parameters) at sufficiently low temperatures can yield a wide panoply of modulated structures, such as cluster crystals, gyroid phasesEdelmann and Roth (2016), crystals of bars, etc.
@@
Figure 8: Vibrational frequency spectra derived from the particles velocity self-correlation at various temperatures and $\rho\sigma^{3}=0.05$. A detail of the low frequency region is presented in the inset in the lower right corner. The cluster vibrational frequency spectrum as computed from the velocity self-correlation of the persistent clusters centers of mass is depicted in the inset in the upper right corner.
@@
Figure 9: Cluster-cluster pair distribution function and intra-cluster pair distribution function (inset) for the lowest temperature state $k_{B}T/\zeta=3$.
@@
Now a few words concerning the single particle dynamics. In what follows, in order to make the results independent of specific details of the model, time and frequency will be reduced with the corresponding molecular time unit, $\tau_{0}$ and its inverse $\tau_{0}^{-1}$ respectively. Here we first look at the individual particle mean square displacement (m.s.d), plotted in the upper graph of Figure 7. It is apparent from the figure that there are clearly several separate regimes. First, there is a short time scale (under 2 $\tau_{0}$, i.e. for our system around 14ps) where the system is in a diffusive regime corresponding to intracluster movements (the mean square displacement is $\lesssim 16\sigma^{2}\approx<R_{g}>^{2}$). Then the system reaches a subdiffusive regime with a decreasing diffusion constant (slope of the curves). This reflects the finite size of the clusters. For much longer times (up to 2.7×$\times 10^{3}\tau_{0}$) the behavior of the mean square displacement is represented in the inset in log-log scale. The low density-high temperature case displays a fully linear behavior in the m.s.d. In this case the regime is consistently diffusive at all times, except for the short lived ballistic region at very short times, hardly visible in the graph. In contrast, the two systems with the largest density display a plateau after the initial diffusive and subdiffusive regimes, which is then followed by a long time slightly diffusive behavior, hardly visible in the case of low T and highest density. This long time behavior stems from the collective cluster movement which is shown in the lower graph of Figure 7. Here we have plotted the mean square displacement of the center of mass of persistent clusters, that is, those that preserve their identity during the length of the production run. At short times one can even appreciate the ballistic regime with a parabolic m.s.d. before the clusters begin to collide. We see that the long time behavior (inset in log-log scale) displays the same trends as the single particle mean square displacement. One can actually measure a tiny diffusivity for the clusters at higher density/low temperature, namely $D/(\sigma^{2}\zeta/m)^{1/2}=0.0003$. If this quantity is rescaled to take into account the average mass of the clusters we would still have a very small diffusivity, $D/(\sigma^{2}\zeta/M)^{1/2}\approx 0.006$. This is an indication that our system is approaching an arrested glassy state, and the dynamic slow down is not only due to the larger mass of the clusters. Taking into account that the source of this quasi-freezing of the cluster positions is mostly induced by the long range of the repulsive interactions (cf the right graph of Fig. 1) one is tempted to identify this state as a precursor of a Wigner glass of clustersRuiz-Franco and Zaccarelli (2021). However, strictly speaking the formation of Wigner crystals and glasses Duang and Guojun (2005) is due to extremely long ranged interactions such a Coulombic ones, in conjunction with higher packing fractions. Following the discussion of Klix and coworkers Klix et al. (2010) and bearing in mind the significant interparticle attraction and low particle packing fraction of our model, we are most likely somewhere near the dynamic transition towards a cluster glassy state. It is worth recalling that despite the very low particle density, as mentioned before, (cf Figure 2) due to the large intercluster repulsion the effective density as inferred from the cluster excluded volume lies in the range $\rho_{C}\lambda_{0}^{3}=0.74\sim 0.82$ for all cases studied here, corresponding to relatively dense fluid states.
These features are further confirmed by an analysis of the velocity self-correlation functions which are illustrated and commented upon in the Supplementary Information (cf. Figure S2). These quantities, once Fourier transformed provide the frequency spectra, which are plotted in Figure 8. We see first that the high values of $Z(0)$ confirm the liquid-like dynamics of the cluster constituent particles. For $k_{B}T/\zeta=3$ and $\rho\sigma^{3}=0.05$, cage effect vibrations are reflected as a tiny maximum (see lower inset of Figure 8) observed at $10^{12}(m\sigma^{2}/\zeta)^{1/2}\nu\approx 0.65$, whereas as T is increased a much wider maximum occurs at lower frequencies $10^{12}(m\sigma^{2}/\zeta)^{1/2}\nu\approx 0.25$. This maximum stems from the coupled particle-cluster dynamics, an effect also seen in the velocity self-correlation function depicted in Figure S2. In the upper inset we have the corresponding vibrational frequency spectrum derived from the persistent cluster velocity self-correlation function. Now the situation is completely different and the low frequency behavior corresponds to a that of a system approaching a glassy state, with very low diffusivities ($Z_{cl}(0)\longrightarrow 0$), in agreement with the long time behavior of the m.s.d. seen in Figure 7. Note that the frequency domain of the cluster dynamics is one order of magnitude smaller than that of single particles, reflecting the much slower dynamics of the clusters, a consequence of their large masses. The intensity of the peaks increases with temperature, as the kinetic energy increase reflects mostly in vibrations with larger amplitude around equilibrium positions, with the maxima shifting to slightly higher energies. This three maxima correspond to three minima (cages) in the cluster velocity self-correlation function (right graph in Figure S2), where also one can see that the cages become narrower and move to shorter times as temperature is increased. This effect is due to the higher vibrational energy of thermal origin present in the high T state, which for $\rho\sigma^{3}=0.05$ preserves some of the characteristics of a quasi-amorphous state. Lowering the density and/or further increasing the temperature above $k_{B}T/\zeta>15$ obviously ends up in the dissociation of the clusters and the complete disappearance of the vibrational structure of the spectrum.
@@
Figure 10: Cluster-cluster structure factor for the lowest temperature state, $k_{B}T/\zeta=3$.. In inset the low-Q behavior is illustrated with using $Q^{2}$ as abscissa.
V Two-particle and intercluster structural correlations. Build-up of effective hyperuniformity